# Material-informed Gaussian Splatting for 3D World Reconstruction in a Digital Twin

### v2

@@ I Introduction
Digital Twins are integral to advancing sensor technologies, facilitating the safe validation of high-risk scenarios and precise control over environmental variables prior to real-world deployment. Modern Advanced Driver Assistance Systems (ADAS) and Artificial Intelligence (AI) systems rely on sophisticated algorithms and complex multimodal sensor setups, often including LiDAR sensors and cameras. Achieving accurate translation of real-world scenarios to simulation—where virtual sensors exhibit identical behavior to their physical counterparts—requires 3D world reconstruction that captures spatial and geometric features to ensure realistic simulation and synthetic sensor data consistent with reality.
Previous LiDAR-based approaches Zaalberg2024 ; Manivasagam have achieved accurate geometric reconstruction for static scenes but lack the texture information naturally captured by cameras, limiting photorealistic rendering capabilities. Moreover, LiDAR struggles with transparent and reflective materials such as glass and metal, which are readily captured in camera images but produce inaccurate point clouds. While incorporating material properties into LiDAR-based reconstruction has been shown to improve sensor simulation accuracy muckenhuber2020automotive , such approaches still require LiDAR hardware and do not address texture acquisition. Combining LiDAR with cameras to capture both geometry and appearance requires complex sensor synchronization and calibration procedures, yet still produces limited texture quality under sparse-view conditions, as traditional reconstruction methods struggle to generate photorealistic results.
Recent advancements in novel view synthesis have introduced 3D Gaussian Splatting kerbl20233dgaussiansplattingrealtime , an efficient technique providing fast, differentiable rendering and high-quality reconstructions from sparse input views. While originally focused on photorealistic rendering, recent works such as MILo guédon2025milomeshintheloopgaussiansplatting have shown that geometric surfaces can be extracted from Gaussian Splatting using only multi-view images, enabling a hybrid representation where Gaussian Splatting provides photorealistic visualization while extracted meshes enable physics-based simulation in modern graphics engines. These camera-only methods provide an alternative to LiDAR hardware, avoiding intricate sensor synchronization and demanding calibration procedures while delivering superior texture quality.
@@ II-A Novel View Synthesis
Traditional 3D reconstruction methods use explicit representations. LiDAR-based methods huang2023neural excel at geometric accuracy but struggle with fine visual details. Camera-based alternatives Schonberger2016 offer richer visual information but produce noisier geometry and face challenges with transparent and reflective surfaces remondino2023critical . Neural Radiance Fields (NeRF) mildenhall2020nerfrepresentingscenesneural enabled photorealistic novel view synthesis from camera-only input but suffers from high computational overhead and lacks discrete surface representation needed for graphics engines.
@@ III-A Monocular Material Extraction
However, material regions often exhibit fragmented spatial distributions and lack the prominent shape cues of semantic objects. To address this, we employ Fast Segment Anything (FastSAM) zhao2023fastsegment to identify all instances within each scene. FastSAM produces pixel-wise boundaries, which we post-process to remove overlaps before overlaying with RMSNet predictions.
For each segmented region, we apply majority voting: we count the material class labels within its boundary and assign the most frequent label to all enclosed pixels. This strategy ensures consistent material assignment while preserving sharp boundaries, as illustrated in Fig. 2. The result is a set of n shape-aware material masks, where each pixel inherits its class according to the FastSAM-derived segmentation.
@@ III-B Large-Scale Gaussian Splatting
Our 3D reconstruction combines photorealistic rendering with explicit geometry through a hybrid approach. We first apply COLMAP Schonberger2016 to generate a sparse point cloud from multi-view RGB images, providing GPS-derived camera poses, pre-calibrated intrinsics, and extrinsics as priors. To avoid artifacts from the moving ego-vehicle, we mask it out in all input images.
From this COLMAP output, we train two complementary models. For photorealistic visualization, we employ H3DGS h3dgs , selected over alternatives such as CityGaussianV2 liu2025citygaussianv2efficientgeometricallyaccurate due to its superior visual quality on street-level autonomous driving datasets. H3DGS constructs a global scaffold and subdivides the scene into chunks. However, its hierarchical level-of-detail representation and anti-aliasing mechanisms are not compatible with the standard format needed for downstream applications. We therefore disable these features and perform a custom merging of the chunks.
For mesh geometry and semantic labeling, we employ MiLO guédon2025milomeshintheloopgaussiansplatting , which provides geometrically accurate mesh extraction required for physics-based simulation. Trained on the same COLMAP sparse point cloud, MiLO produces both a 3D Gaussian model and an explicit mesh. The MiLO Gaussian model initializes our semantic Gaussian Splatting for material projection (Sec. III-C), while the mesh provides the geometric basis for physics-based simulation (Sec. III-E). Both H3DGS and MiLO reconstructions share the same coordinate system, ensuring alignment between appearance and geometry.
@@ III-C Per-Pixel Material Projection
This module projects the 2D material masks from Sec. III-A onto the MiLO Gaussian representation, then transfers the labels to the mesh. Unlike previous pipelines relying on camera-to-LiDAR transformations, our Gaussian model is natively aligned in world coordinates. We employ differentiable rasterization zwicker2001ewa to render Gaussians to 2D, maintaining correspondence between 3D Gaussian positions and 2D pixel coordinates. When multiple Gaussians project to the same pixel, depth sorting during alpha blending assigns the label of the closest Gaussian.
@@ III-D Physics-Based Material Assignment
We assign physics-based material properties to the mesh using the Principled BSDF (Bidirectional Scattering Distribution Function) shader following the Disney BSDF standard burley2012physically , ensuring compatibility with graphics engines such as Blender, Unreal Engine, and simulation platforms like Simcenter Prescan. The Principled BSDF is parameterized by physically meaningful properties including base color, roughness, metallic factor, and transparency, ensuring physically accurate and consistent material representation.
Materials are sourced from the Matsynth dataset Vecchio_2024 , which provides over 4,000 physically-based materials across 14 categories. These datasets provide PBR (Physically-Based Rendering) texture maps including base color, normal maps, roughness, and metallic maps. For each material class from Sec. III-A, we select representative PBR textures and apply them to the corresponding mesh triangles, producing a fully textured mesh directly usable by standard graphics engines and physics-based sensor simulation frameworks.
@@ IV-A Dataset
We collected data using an instrumented test vehicle equipped with an Ouster OS1-128 LiDAR, six surround-view cameras, and RTK-GPS, all hardware-synchronized at 20Hz. Five static road scenes were captured in an urban environment, each approximately eight seconds long, recorded from a moving ego-vehicle.
@@ IV-C Implementation Details
Scene reconstruction: We train all Gaussian Splatting methods using their default hyperparameters. Material assignment follows the pipeline described in Sec. III-AIII-D, projecting semantic labels onto the reconstructed geometry and mapping them to PBR material properties.
@@ IV-E Visual Quality Evaluation
Quantitative analysis. Table I presents PSNR, SSIM, and LPIPS metrics across all methods and scenes. Our adapted model achieves lower PSNR (18.27 vs. 20.07–21.22) and SSIM (0.41 vs. 0.56–0.58) compared to baselines. However, it achieves the best average LPIPS score (0.44 vs. 0.46–0.53), outperforming all methods on four out of five scenes. This is particularly significant since LPIPS correlates strongly with human perceptual similarity zhang2018unreasonableeffectivenessdeepfeatures , suggesting that our reconstruction maintains perceptual quality despite lower pixel-level accuracy. The lower PSNR and SSIM values potentially stem from our architectural modifications: disabling H3DGS’s hierarchical level-of-detail representation and anti-aliasing mechanisms (Sec. III-B) trades pixel-level accuracy for compatibility with downstream simulation engines. However, the superior LPIPS score suggests that these modifications preserve perceptual quality.
@@ IV-F LiDAR Simulation Accuracy
Qualitative analysis. Figure 4 presents a visual comparison of LiDAR reflectivity predictions between our camera-only method, the LiDAR-camera fusion baseline, and real-world ground truth. Both methods exhibit reflectivity errors, with vegetation consistently showing higher predicted reflectivity than ground truth—likely due to the diversity of vegetation forms and our simplified PBR material parameterization. The camera-only reconstruction exhibits more geometric noise than the LiDAR-based approach, as expected from camera-based methods. Despite these differences, both simulation approaches produce comparable reflectivity patterns, validating our camera-only pipeline for sensor simulation applications.
@@
Quantitative analysis. Table II and Figure 5 present reflectivity error metrics across all five scenes. Our camera-only method achieves comparable reflectivity accuracy to the LiDAR-camera fusion baseline, with mean absolute error (MAE: 10.05 vs. 10.14) and median error (6.48 vs. 6.78) showing minimal differences between the two approaches. This demonstrates that camera-only reconstruction can achieve sensor simulation accuracy comparable to LiDAR-based methods without requiring LiDAR hardware or complex sensor synchronization. The higher within-scene variability in camera-only results (visible in Figure 5) reflects the increased geometric noise inherent to camera-based reconstruction, which lacks the direct depth measurements provided by LiDAR.
@@ V Conclusion
Future work. Future directions include improving geometric reconstruction quality from camera inputs, extending to dynamic scenes, and exploring learned material representations to reduce dependency on predefined PBR databases.

### v3

@@ I Introduction
Digital Twins advance sensor technologies by enabling safe validation of high-risk scenarios and precise environmental control prior to real-world deployment. Modern Advanced Driver Assistance Systems (ADAS) and Artificial Intelligence (AI) systems rely on sophisticated algorithms and complex multimodal sensor setups, often including LiDAR sensors and cameras. Achieving accurate translation of real-world scenarios to simulation—where virtual sensors exhibit identical behavior to their physical counterparts—requires 3D world reconstruction that captures spatial and geometric features to ensure realistic simulation and synthetic sensor data consistent with reality.
Previous LiDAR-based approaches Zaalberg2024 ; Manivasagam achieve accurate geometry but lack camera-captured texture information. Moreover, LiDAR struggles with transparent and reflective materials like glass and metal, readily captured in camera images. While incorporating material properties into LiDAR-based reconstruction has been shown to improve sensor simulation accuracy muckenhuber2020automotive , such approaches still require LiDAR hardware and do not address texture acquisition. Combining LiDAR with cameras to capture both geometry and appearance requires complex sensor synchronization and calibration procedures, yet still produces limited texture quality under sparse-view conditions, as traditional reconstruction methods struggle to generate photorealistic results.
Recent advancements in novel view synthesis have introduced 3D Gaussian Splatting kerbl20233dgaussiansplattingrealtime , an efficient technique providing fast, differentiable rendering and high-quality reconstructions from sparse input views. Recent works like MILo guédon2025milomeshintheloopgaussiansplatting extract geometric surfaces from Gaussian Splatting using only multi-view images, enabling hybrid representations combining photorealistic visualization with physics-based simulation in graphics engines. These camera-only methods provide an alternative to LiDAR hardware, avoiding intricate sensor synchronization and demanding calibration procedures while delivering superior texture quality.
@@ II-A Novel View Synthesis
Traditional 3D reconstruction methods use explicit representations. LiDAR-based methods huang2023neural excel at geometric accuracy but struggle with fine visual details. Camera-based alternatives Schonberger2016 offer richer visual information but produce noisier geometry and face challenges with transparent and reflective surfaces remondino2023critical . Neural Radiance Fields (NeRF) mildenhall2020nerfrepresentingscenesneural enabled photorealistic novel view synthesis from camera-only input but suffers from high computational overhead and lacks discrete surfaces needed for graphics engines.
@@ III-A Monocular Material Extraction
Material regions often exhibit fragmented distributions lacking prominent shape cues. We employ FastSAM zhao2023fastsegment to identify instances, producing pixel-wise boundaries that we post-process to remove overlaps before overlaying with RMSNet predictions.
For each region, we apply majority voting: counting material class labels within its boundary and assigning the most frequent label to all enclosed pixels. This strategy ensures consistent material assignment while preserving sharp boundaries, as illustrated in Fig. 2. The result is a set of n shape-aware material masks, where each pixel inherits its class according to the FastSAM-derived segmentation.
@@ III-B Large-Scale Gaussian Splatting
Our 3D reconstruction combines photorealistic rendering with explicit geometry through a hybrid approach. We first apply COLMAP Schonberger2016 to generate a sparse point cloud from multi-view RGB images, providing GPS-derived camera poses, pre-calibrated intrinsics, and extrinsics as priors. We mask the moving ego-vehicle in all input images to avoid artifacts.
From the COLMAP output, we train two complementary models. For photorealistic visualization, we use H3DGS h3dgs , selected for its superior visual quality on autonomous driving datasets over alternatives like CityGaussianV2 liu2025citygaussianv2efficientgeometricallyaccurate . H3DGS constructs a global scaffold and subdivides the scene into chunks. However, H3DGS’s hierarchical LoD representation uses a proprietary .hier format, and its anti-aliasing mechanism yu2023mip is not supported by standard 3DGS rendering tools that expect the conventional .ply format. To maximize compatibility with existing 3DGS viewers and downstream processing tools, we therefore disable these features and perform a custom merging of the chunks.
For mesh geometry and semantic labeling, we employ MiLO guédon2025milomeshintheloopgaussiansplatting , which provides geometrically accurate mesh extraction required for physics-based simulation. Trained on the same COLMAP sparse point cloud, MiLO produces both a 3D Gaussian model and an explicit mesh. The MiLO Gaussian model initializes our semantic Gaussian Splatting for material projection (Sec. III-C), while the mesh provides the geometric basis for physics-based simulation (Sec. III-E). Both H3DGS and MiLO are trained with monocular depth supervision depth_anything_v2 following their respective default configurations, and both reconstructions share the same coordinate system, ensuring alignment between appearance and geometry.
@@ III-C Per-Pixel Material Projection
This module projects the 2D material masks from Sec. III-A onto the MiLO Gaussian representation, then transfers the labels to the mesh. Unlike previous pipelines relying on camera-to-LiDAR transformations, our Gaussian model is natively aligned in world coordinates. We employ differentiable rasterization to render Gaussians to 2D, maintaining correspondence between 3D Gaussian positions and 2D pixel coordinates. When multiple Gaussians project to the same pixel, depth sorting during alpha blending assigns the label of the closest Gaussian.
@@ III-D Physics-Based Material Assignment
We assign physics-based material properties to the mesh using the Principled BSDF shader following the Disney BSDF standard burley2012physically , ensuring compatibility with graphics engines such as Blender, Unreal Engine, and simulation platforms like Simcenter Prescan. The Principled BSDF is parameterized by physically meaningful properties including base color, specular, metallic factor, roughness, opacity, and surface normals.
From the 20 material classes detected by RMSNet (Sec. III-A), we selected a subset commonly encountered in urban driving environments: glass, brick and ceramics, concrete, asphalt, vegetation (e.g., grass and leaves), metals, plastics, gravel, tree trunks, and rubber. For each category, PBR textures were sourced from the Matsynth dataset Vecchio_2024 , which provides over 4,000 physically-based materials across 14 categories, and matched to Prescan’s laboratory-tested material database to ensure validated reflectivity properties for LiDAR simulation. Within each category, textures were balanced across key parameters (base color, roughness, metallic factor, clear coat). For each material class from Sec. III-A, we apply the selected PBR textures to the corresponding mesh triangles, producing a fully textured mesh directly usable by standard graphics engines and physics-based sensor simulation frameworks.
@@ IV-A Dataset
We collected data using an instrumented test vehicle equipped with an Ouster OS1-128 LiDAR ousterOS1Specs (128 channels, 20 Hz, 100 Hz IMU), six surround-view cameras with Sony IMX490 sensors (2880×1860 resolution, one rear with fisheye lens for 360° coverage), and a Septentrio AsteRx GNSS/RTK system (20 Hz). The LiDAR provides hardware triggers to synchronize all cameras at 20 Hz. Five static road scenes were captured near our company facilities in Leuven, Belgium, each approximately eight seconds at 20-30 km/h. Data collection on company-controlled roads avoided GDPR compliance issues and ensured controlled validation conditions.
We focus on static scene reconstruction as this aligns with Digital Twin simulation workflows. Modern simulation platforms (e.g., Simcenter Prescan, CARLA) provide realistic dynamic actor models with controllable behaviors, making the challenge the creation of photorealistic static environments (roads, buildings, vegetation, infrastructure). Our method automates this process, providing realistic backgrounds for simulation scenarios. Dynamic actors in our recordings are masked out during reconstruction. While our detection pipeline enables trajectory replication for re-insertion as simulation assets, this falls outside the scope of this paper.
Scenes primarily consist of company parking lots and nearby access roads, capturing typical automotive testing environments with diverse material challenges including asphalt, concrete, glass facades, metallic surfaces, extensive vegetation (trees, hedges, grass), road markings, and signage. Recordings under varying weather conditions (clear to partly cloudy) capture natural lighting variations. We reserve 50 novel viewpoints per scene for evaluation.
@@ IV-C Implementation Details
Scene reconstruction: We train all Gaussian Splatting methods using their default hyperparameters as specified in their respective repositories. For both H3DGS and MiLO, we enable monocular depth supervision using Depth Anything V2 depth_anything_v2 . Material assignment follows the pipeline described in Sec. III-AIII-D, projecting semantic labels onto the reconstructed geometry and mapping them to PBR material properties.
@@ IV-E Visual Quality Evaluation
Quantitative analysis. Table I presents PSNR, SSIM, and LPIPS metrics across all methods and scenes. Our adapted H3DGS model is used solely for photorealistic visualization, while MiLO provides geometry for simulation (Sec. III-B). The model achieves lower PSNR (18.27 vs. 20.07–21.22) and SSIM (0.41 vs. 0.56–0.58) compared to baselines, but the best LPIPS score (0.44 vs. 0.46–0.53), outperforming all methods on four out of five scenes. Since LPIPS correlates strongly with human perceptual similarity zhang2018unreasonableeffectivenessdeepfeatures , this suggests our reconstruction maintains perceptual quality despite lower pixel-level accuracy.
The degradation stems from disabling H3DGS’s hierarchical optimization and anti-aliasing mechanism yu2023mip . As noted by the H3DGS authors, hierarchies operating at different scales require correct anti-aliasing h3dgs . Anti-aliasing alone improves 3DGS quality by +4.56 dB PSNR and +0.070 SSIM yu2023mip ; our observed degradation (-1.80 dB PSNR, -0.15 SSIM) is consistent with these modifications. The superior LPIPS score confirms these changes preserve perceptual quality while ensuring compatibility with standard rendering tools.
@@ IV-F LiDAR Simulation Accuracy
Qualitative analysis. Figure 4 presents a visual comparison of LiDAR reflectivity predictions between our camera-only method, the LiDAR-camera fusion baseline, and real-world ground truth. Both methods consistently overestimate vegetation reflectivity due to: (1) temporal inconsistency from wind-induced vegetation movement violating static scene assumptions, (2) high intra-class material variance (leaves, bark, grass share one label but have different reflectivities), (3) idealized PBR materials not capturing weathered outdoor conditions, and (4) geometric complexity affecting surface normal estimation. These limitations impact both reconstruction approaches equally, indicating bottlenecks in the static scene assumption and material parameterization rather than reconstruction modality. Despite higher geometric noise in the camera-only method, both approaches produce comparable reflectivity patterns, validating our pipeline for sensor simulation.
@@
Quantitative analysis. Table II and Figure 5 present reflectivity error metrics across all five scenes. Our camera-only method achieves comparable reflectivity accuracy to the LiDAR-camera fusion baseline, with mean absolute error (MAE: 10.05 vs. 10.14) and median error (6.48 vs. 6.78) showing minimal differences between the two approaches. This demonstrates that camera-only reconstruction can achieve sensor simulation accuracy comparable to LiDAR-based methods without requiring LiDAR hardware or complex sensor synchronization. The higher within-scene variability in camera-only results (visible in Figure 5) reflects the increased geometric noise inherent to MiLO’s camera-based mesh extraction, which lacks the direct depth measurements provided by LiDAR. However, the comparable MAE and median errors demonstrate that this geometric uncertainty does not significantly degrade reflectivity simulation accuracy.
@@ V Conclusion
Future work. Future directions include improving geometric reconstruction quality from camera inputs, extending to dynamic scenes, and exploring learned material representations to reduce dependency on predefined PBR databases. Specific improvements for vegetation modeling could include: (1) environment-specific material calibration capturing weathering and seasonal effects, (2) hierarchical material classification distinguishing foliage types, bark, and ground cover, and (3) refined geometric alignment between multi-modal reconstructions to improve surface normal accuracy for reflectivity calculations. We have developed an internal tool to convert our data to the nuScenes format, enabling future benchmarking against public datasets to demonstrate broader generalization across diverse urban environments.